the sum of three consecutive odd integers is 201. Find the integers.
Found 2 solutions by 13114, ganesh:
Answer by 13114(5)

Let x, x + 2, x + 4 be the three consecutive odd integers.
Then x + x + 2 + x + 4 = 201
3x + 6 = 201
3x = 195
x = 65
The three integers are 65, 67,and 69.

Answer by ganesh(20)

Let x be an odd integers.
When you add 2 with it, you will get the next odd integers.
So, x + 2 is the next odd integers.
Further if you add 2, you will get the next odd integers.
So, x + 2 + 2 = x + 4 is the next odd integers.
So, x, x + 2 and x + 4 are three consecutive odd integers.
It is given that the sum of these three numbers is 201.
Therefore, x + x + 2 + x + 4 = 20.
Or, 3x + 6 = 201
Or, 3x = 195
Or, x = 195/3
Or, x = 65
So, x + 2 = 67, x + 4 = 69.
Therefore, 65, 67 and 69 are the required numbers.